OpenStudy (anonymous):

A pilot flies a plane south and then 600 miles west where she land the plane. How far south did the pilot fly the plane if she lands 610 miles from her starting point.

OpenStudy (anonymous):

@mathslover @YanaSidlinskiy

OpenStudy (yanasidlinskiy):



OpenStudy (yanasidlinskiy):

\[a^2 +b^2 =c^2\] \[x^2+600^2=610^2\] \[x^2+360000=372100 \] \[x^2=12100\] \[x=110 miles\]

OpenStudy (anonymous):

Thank you!

OpenStudy (yanasidlinskiy):

No problem!!:) It is Simple and easy!! Any doubts/questions??

OpenStudy (anonymous):

Nope! Can I ask you one more question?

OpenStudy (yanasidlinskiy):

Sure..

OpenStudy (anonymous):

Simplify. \[4\sqrt{6}+\sqrt{54}\]

OpenStudy (yanasidlinskiy):

That one I'm NOT good at. You might want to get help from someone smarter than me. Like @mathslover .

mathslover (mathslover):

Yep sure, I will try my best to help @pickles112233 @pickles112233 we can write 54 as 9 * 6 , right?

OpenStudy (anonymous):

Yeah.

mathslover (mathslover):

Okay, so I can write the expression as : \(4\sqrt{6} + \sqrt{9 \times 6}\) Now, as 9 = 3^2 so, 3 will come out

mathslover (mathslover):

So, we get : \(4\sqrt{6} + 3\sqrt{6}\)

mathslover (mathslover):

Take \(\sqrt{6}\) common, we get : \(\sqrt{6}(4+3) = 7\sqrt{6}\)

OpenStudy (anonymous):

Thank you so much!

mathslover (mathslover):

You're welcome and Welcome to OpenStudy! :)